1a. Calculate the pH of a 1.35 × 10−4M HNO3 solution. Report your answer to TWO...

1a.

Calculate the pH of a 1.35 × 10−4M HNO3 solution.

Report your answer to TWO places past the decimal.

1b.

Calculate the pH of a 3.25 × 10−6M HClO3 solution.

Report your answer to TWO places past the decimal.

1c

Calculate the pH of a 0.00852 M RbOH solution.

Report your answer to TWO places past the decimal.

1d.

Calculate the pH of a solution that contains 0.0727 g of Ca(OH)2 in 1.83 L of water.

Report your answer to TWO places past the decimal.
Answer #1

1/2008 D 6 (HNO3) = [H +) = 1.35 x1024 pH = -log [ht] pH = -log (1.35 810-4) = -lug 1:35 - log 104 -pH = 4 - 0 13 = 3177 pH =